A rectangle's width is 1 meter more than half of its length. Find the dimensions if the perimeter measures 53 meters.
Answer by stanbon(26274)

A rectangle's width is 1 meter more than half of its length. Find the dimensions if the perimeter measures 53 meters.
-------------
Let length = L
The width = (1/2)L+1
------------------------------
Perimeter = 2(W + L)
53 = 2((1/2L+1 + L)
53 = 2((3/2)L + 1)
52 = 3L + 2
3L = 51
L = 17 meters (length)
W = (1/2)L + 1 = 8.5+1 = 9.5 meters (width)
=============================================
Cheers,
Stan H.